### v1

# Graph-Based Fault Diagnosis for Rotating Machinery: Adaptive Segmentation and Structural Feature Integration

Abstract
This paper proposes a novel graph-based framework for robust and interpretable multiclass fault diagnosis in rotating machinery. The method integrates entropy-optimized signal segmentation, time–frequency feature extraction, and graph-theoretic modeling to transform vibration signals into structured representations suitable for classification. Graph metrics—such as average shortest path length, modularity, and spectral gap—are computed and combined with local features to capture both global and segment-level fault characteristics. Evaluated on two benchmark datasets, the CWRU bearing dataset (under 0–3 HP loads) and the SU gearbox and bearing datasets (under different speed–load configurations), the proposed method achieves high diagnostic accuracy. Classification scores reach up to 99.8% accuracy on Case Western Reserve University (CWRU) and 100% accuracy on the Southeast University datasets using a logistic regression classifier. Furthermore, the model exhibits strong noise resilience, maintaining over 95.4% accuracy at high noise levels (standard deviation = 0.5), and demonstrates excellent cross-domain transferability with up to 99.7% F1-score in load-transfer scenarios. Compared to traditional techniques, this approach requires no deep learning architecture, enabling lower complexity while ensuring interpretability. The results confirm the method’s scalability, reliability, and potential for real-time deployment in industrial diagnostics.
keywords:
Bearing fault diagnosis, entropy-based segmentation, feature noise robustness, graph-based learning, graph-theoretic features, multiclass classification, transfer learning, vibration signal analysis.
journal: Nuclear Physics B
[…]
[label1]organization=Department of Computer Science and Engineering, DUIET, Dibrugarh University, addressline=Dibrugarh, postcode=786004, state=Assam, country=India
1 Introduction
Rotating machinery, such as motors, gearboxes, and bearings, is fundamental to industries including manufacturing, energy, transportation, and aerospace. The reliability of these systems depends on early fault detection in critical components like bearings and gearboxes, which are highly susceptible to failure due to mechanical stress (Lei et al., 2013). Undetected faults can cause costly downtime, maintenance expenses, or safety hazards, necessitating accurate and efficient fault diagnosis systems (Jardine et al., 2006).
Traditional fault diagnosis methods rely on statistical and spectral analysis of vibration signals, using fixed-window segmentation, handcrafted feature extraction (e.g., time- or frequency-domain metrics), and classifiers like support vector machines (Widodo and Yang, 2007). While effective in controlled settings, these approaches have limitations. Fixed segmentation struggles to adapt to dynamic signal variations across operating conditions, leading to suboptimal feature representations (Rauber et al., 2014). Handcrafted features often fail to capture complex or subtle fault patterns, especially in noisy environments (Lei et al., 2020). Additionally, these methods assume consistent training and testing conditions, which is rare in practice, resulting in poor generalization across varying loads, fault severities, or noise levels (Zhang et al., 2017).
Recent advances in deep learning have enhanced fault diagnosis capabilities. Convolutional neural networks (CNNs) (Wen et al., 2017), recurrent neural networks (RNNs) (Zhu et al., 2023), and attention-based models (Zhang et al., 2020a) have achieved strong results on benchmark datasets like the Case Western Reserve University (CWRU) bearing dataset (Center, 2014) and the Southeast University (SU) gearbox dataset(Wu, 2020). However, these methods require large labeled datasets, incur high computational costs, and lack interpretability due to their black-box nature (Zhang et al., 2020b). Their performance also degrades in cross-domain or cross-load scenarios, where training and deployment conditions differ (Li et al., 2022).
A critical gap in existing approaches is the underuse of structural information within vibration signals. These signals often exhibit temporal relationships reflecting physical dependencies, yet conventional methods process segments independently, overlooking inter-segment patterns (Yang et al., 2021). Graph-based representations offer a promising solution, enabling the modeling of structural relationships to enhance fault detection.
To address these challenges, this paper proposes a novel graph-based framework for multiclass fault diagnosis in rotating machinery. Vibration signals are segmented adaptively using entropy-based optimization of window and step sizes, ensuring robust feature capture across diverse operating conditions. Each segment is represented by a comprehensive feature set, including time-domain statistics, frequency-domain metrics (e.g., spectral density, envelope analysis), and Teager-Kaiser energy (Welch, 1967). These features are clustered to form compact representations, from which k-nearest neighbor (kNN) graphs are constructed. Graph-theoretic metrics—such as average shortest path length, modularity, and spectral gap—encode fault-induced structural variations, capturing disruptions in signal regularity and community structure (Newman, 2018).
The framework integrates global graph metrics with local signal features, enabling a dual-level representation that distinguishes subtle fault types and enhances noise robustness. A multiclass logistic regression classifier, leveraging these high-quality features, achieves state-of-the-art performance with computational efficiency and interpretability, unlike complex deep learning models.
The objectives of this study are as follows:
  • 1.
    To develop an adaptive segmentation technique based on entropy-driven optimization.
  • 2.
    To construct graph representations from feature-rich segments of vibration signals.
  • 3.
    To extract graph-theoretic metrics that effectively capture fault-related characteristics.
  • 4.
    To evaluate the proposed framework on the CWRU and SU datasets across varying load conditions, fault types, noise levels, and cross-domain (transfer learning) scenarios.
Extensive experiments demonstrate that the proposed method achieves high accuracy, robust generalization, and resilience to noise, outperforming traditional and deep learning approaches. Its efficiency and interpretability make it well-suited for industrial applications. These results highlight the potential of graph-based learning to advance intelligent fault diagnosis, offering a scalable and practical solution for real-world challenges.
2 Literature Review
Traditional fault diagnosis methods rely on handcrafted features extracted from vibration signals, typically in the time or frequency domain, followed by classifiers like support vector machines (SVMs) (Widodo and Yang, 2007). Time-domain features (e.g., mean, RMS, skewness (Joanes and Gill, 1998)) and frequency-domain features (e.g., FFT (Welch, 1967), spectral envelope (Zheng et al., 2022)) capture fault signatures, while techniques like empirical mode decomposition (EMD) (Lei et al., 2013) and Teager-Kaiser energy (Kaiser, 1990) enhance signal decomposition. These methods achieve high accuracy in controlled settings, such as 95–98% on the CWRU dataset (Zhang et al., 2015), but struggle with dynamic operating conditions and noise (Rauber et al., 2014). Fixed segmentation and manual feature selection limit adaptability, requiring domain expertise (Lei et al., 2020).
Machine learning has improved fault diagnosis by automating feature selection and classification. Ensemble methods (Zhang et al., 2015) and transfer learning (Zhang et al., 2017) enhance robustness across operating conditions. For instance, Zhang et al. (Zhang et al., 2017) used neural networks for bearing fault diagnosis under varying loads, achieving 90–95% accuracy on CWRU cross-load scenarios. However, these methods still rely on handcrafted features, limiting their ability to capture complex patterns in noisy or high-dimensional data (Lei et al., 2020). Classifiers like k-nearest neighbors (kNN) (Cover and Hart, 1967) and SVMs (Widodo and Yang, 2007) are computationally efficient but lack the generalization of deep learning for large datasets.
Deep learning has revolutionized fault diagnosis by learning hierarchical features from raw vibration signals (Zhao et al., 2019). Convolutional neural networks (CNNs) (Wen et al., 2017; Niyongabo et al., 2022; Xu et al., 2025), long short-term memory (LSTM) networks (Zhang et al., 2020a), and attention-based models (Li et al., 2024) achieve high accuracies (98–99.9%) on benchmarks like CWRU and SU (Center, 2014; Wu, 2020). For example, Shao et al. (Shao et al., 2018) reported 99.5% accuracy using deep transfer learning on SU. Recent advancements include multi-source fusion (Makrouf et al., 2023), sparse CNNs (Xu et al., 2025), and transformer-based models (Zhou et al., 2024), improving efficiency and robustness. However, deep learning models require large labeled datasets, incur high computational costs (e.g., 536 s inference for DL-CNN (Shao et al., 2018)), and lack interpretability (Zhang et al., 2020b). Performance drops in cross-domain scenarios (e.g., 90–95% F1-score on cross-load CWRU (Li et al., 2022)) highlight generalization challenges (Zhu et al., 2023).
Graph-based methods offer a promising approach for modeling structural relationships in vibration signals, advancing fault diagnosis in rotating machinery (Yang et al., 2021; Singh, 2025). Yang et al. (Yang et al., 2021) employed spatial-temporal graphs to extract features, achieving 97% accuracy on the CWRU dataset, while Singh (Singh, 2025) integrated graph autoencoders with transformers for robust fault detection. These methods leverage graph-theoretic metrics, such as modularity (Newman, 2006), spectral gap (Cheeger, 1970; Chung, 1996), and small-world properties (Watts and Strogatz, 1998), to capture fault-induced disruptions in signal structure. While effective at modeling inter-segment dependencies, they are computationally intensive and limited in extracting frequency-domain features, as they prioritize structural relationships over spectral analysis (Yang et al., 2021; Lei et al., 2020). Few studies combine local signal characteristics with global graph metrics, a gap addressed by the proposed framework, which integrates both for enhanced accuracy and interpretability.
The CWRU bearing dataset (Center, 2014) and SU gearbox dataset (Wu, 2020) are widely used for benchmarking. CWRU includes vibration data under varying loads (0–3 HP) and fault types, while SU covers gearbox faults under diverse conditions (Shao et al., 2018). Performance metrics like accuracy, F1-score, and inference time are common, with statistical tests (e.g., ANOVA (Fisher, 1925), Tukey HSD (Tukey, 1949)) used to evaluate cross-load robustness (Zhang et al., 2017). Recent studies report accuracies of 98–99.9% on CWRU and SU, but cross-load performance remains a challenge (e.g., 90–95% F1-score for deep learning (Li et al., 2022)).
Despite significant advancements in vibration-based fault diagnosis of rotating machinery, several critical gaps and challenges persist across traditional, machine learning, deep learning, and graph-based approaches. Traditional and machine learning methods, reliant on manual feature engineering, lack adaptability to dynamic operating conditions and require extensive domain expertise to design effective features, limiting their performance in noisy or variable environments (Lei et al., 2020; Rauber et al., 2014). Deep learning models, while capable of automatically learning complex patterns, demand large labeled datasets and incur high computational costs, often rendering them impractical for real-time industrial applications (Zhu et al., 2023; Shao et al., 2018). Moreover, their black-box nature hampers interpretability, with post-hoc methods like Grad-CAM providing limited insight into fault mechanisms (Zhang et al., 2020b). Both traditional and deep learning approaches struggle with cross-domain generalization, as performance degrades significantly in cross-load or cross-domain scenarios due to domain shifts (Zhang et al., 2017; Li et al., 2022). Graph-based methods, although promising for modeling structural relationships in vibration signals, are underutilized and computationally intensive, often failing to integrate frequency-domain insights critical for fault detection (Yang et al., 2021; Singh, 2025). Additionally, the underuse of inter-segment relationships in vibration signals across all methods misses opportunities to capture fault-induced structural patterns, highlighting the need for approaches that combine adaptive feature extraction, structural modeling, and robust generalization (Lei et al., 2020; Yang et al., 2021).
3 Methodology
We propose a methodology that begins by extracting time- and frequency-domain features $\mathbf{F}_{i}$ from a multichannel vibrational signal $\mathbf{S}$. The signal is segmented, and clustered segments are used to construct a K-Nearest Neighbors (KNN) graph $\mathcal{G}$. From this graph, we compute key graph-theoretic features, including average shortest path length $L_{\text{avg}}$, modularity $Q_{\text{mod}}$, and spectral gap $\Delta_{\text{spec}}$. These features serve as input to a multiclass logistic regression model designed for fault detection. An overview of the proposed pipeline is illustrated in Figure 1, outlining the progression from raw signal input to fault classification via feature extraction, graph construction, graph-based feature computation, and model training. In the following subsection, we detail each component of the proposed approach.
Figure 1: Overview of the methodology for multiclass fault detection.
Table 1 summarizes the key symbols used in this work, focusing on essential variables for brevity.
Table 1: Summary of Key Symbols
Symbol Description
$\mathbf{S}$ Original multi $ch$-channel signal ($\mathbb{R}^{N\times ch}$)
$\omega^{*}$ Optimal window size for segmentation
$\sigma^{*}$ Optimal step size for segmentation
$n$ Number of segments ($\lfloor(N-\omega^{*})/\sigma^{*}\rfloor$)
$\mathbf{F}_{i}$ Feature vector for the $i$-th segment ($\mathbb{R}^{d}$)
$d$ Dimension of feature vectors
$K_{c}$ Number of clusters in MiniBatch K-Means
$k_{\text{NN}}$ Number of nearest neighbors in KNN graph
$\tau$ Distance threshold for KNN graph edges
$\mathcal{G}$ Set of subgraphs $\{\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{m_{G}}\}$
$m_{G}$ Number of subgraphs
$N_{\text{max}}$ Maximum number of nodes per subgraph
$w_{ij}$ Edge weight between nodes $i$ and $j$
$N_{\text{total}}$ Total number of nodes across all subgraphs ($\approx n$)
$L_{\text{avg}}$ System-wide average shortest path length
$Q_{\text{mod}}$ System-wide average modularity
$\Delta_{\text{spec}}$ System-wide average spectral gap
$\ell_{v}$ Multiclass fault label of node $v$ ($\{0,1,\ldots,K_{F}-1\}$)
$K_{F}$ Number of fault classes
$\mathbf{z}_{v}$ Combined feature vector for node $v$ ($\mathbb{R}^{d+3}$)
$\mathbf{Z}$ Feature matrix ($\mathbb{R}^{N_{\text{total}}\times(d+3)}$)
^$\hat{\ell}_{v}$ Predicted class label for node $v$
3.1 Optimizing Segmentation Process
In the vibrational signal fault detection pipeline, we propose a methodology to optimize window parameters for signal segmentation. This approach combines multiple entropy measures with a systematic parameter search to determine the optimal window size ($\omega^{*}$) and step size ($\sigma^{*}$). By leveraging diverse entropy metrics, the method captures complementary signal characteristics, enabling effective segmentation that enhances fault detection accuracy.
We include the step size ($\sigma$) to control the degree of overlap between consecutive segments during signal segmentation. While the window size ($\omega$) determines the duration of each segment, the step size governs how far the window advances across the signal. Choosing an appropriate step size is essential for balancing temporal resolution and redundancy—smaller values of $\sigma$ yield higher overlap and finer temporal detail but increase computational cost and potential data redundancy. Conversely, larger step sizes reduce overlap, speeding up computation but potentially missing transient features. Therefore, optimizing both $\omega$ and $\sigma$ ensures effective signal representation for entropy-based feature extraction.
We begin by calculating Shannon entropy as the foundational metric for a discrete probability distribution $P=\{p_{1},p_{2},\dots,p_{n}\}$, expressed as:
$H(P)=-\sum_{i=1}^{n}p_{i}\log(p_{i})$ (1)
where $p_{i}$ represents the probability of the $i$-th outcome (Shannon, 1948).
For a continuous-time signal $x(t)$, we derive the probability distribution $P$ from amplitude histograms of windowed segments $x_{w}(t)$ of duration $\omega$. The discrete representation of each segment is
$\mathbf{x}_{w}=[x(t_{0}),x(t_{0}+\Delta t),\dots,x(t_{0}+\omega)]\in\mathbb{R}^{\omega},$
where $\Delta t$ denotes the sampling interval and histograms are constructed with a bin size $\beta$, dynamically calculated using the Freedman-Diaconis rule:
$\beta=\lceil 2\cdot\text{IQR}(\mathbf{x}_{w})\cdot\omega^{-1/3}/(\max(\mathbf{x}_{w})-\min(\mathbf{x}_{w}))\rceil$
where $\text{IQR}(\mathbf{x}_{w})$ denotes the interquartile range (Freedman and Diaconis, 1981). Each bin’s probability is estimated as:
$p_{i}=n_{i}/(\omega+\epsilon),$
where $n_{i}$ is the bin count and $\epsilon=10^{-10}$ is a regularization constant to avoid numerical instability.
We computed four specific entropy measures to comprehensively characterize each windowed segment:
  1. 1.
    Amplitude Entropy ($H_{a}$): Calculated directly from the raw signal segment $\mathbf{x}_{w}$ using the histogram described above:
    $H_{a}=-\sum_{i=1}^{\beta}p_{i}\log(p_{i})$
  2. 2.
    Envelope Entropy ($H_{e}$): Captures amplitude variations by applying the Shannon entropy formulation to the analytic signal envelope $|H\{x_{w}(t)\}|$, where $x_{w}(t)$ is the windowed segment of duration $\omega$. The envelope is obtained via the Hilbert transform, defined as
[…]
    where $x_{w}(t)$ is the original real-valued signal within the window and $j=\sqrt{-1}$ is the imaginary unit. The envelope is:
    $|H\{x_{w}(t)\}|=\sqrt{x_{w}(t)^{2}+\hat{x}_{w}(t)^{2}}$
    where $\hat{x}_{w}(t)$ denotes the Hilbert transform of $x_{w}(t)$. The envelope entropy is computed from a histogram of $|H\{x_{w}(t)\}|$ with $\beta$ bins, as follows:
    $H_{e}=-\sum_{i=1}^{\beta}p_{i}\log(p_{i})$
    where $p_{i}$ is the probability of the envelope amplitude being in the ( i )th bin.
  3. 3.
    Spectral Entropy ($H_{s}$): Estimated in the frequency domain using Welch’s method to calculate the power spectral density (PSD) of the windowed segments $\mathbf{x}_{w}$ (Welch, 1967). For a discrete signal segment, we compute
    $S_{xx}(f)=\frac{1}{K}\sum_{i=1}^{K}|X_{i}(f)|^{2}$
    where $X_{i}(f)$ are the Fourier transforms of $K$ overlapping segments of $\mathbf{x}_{w}$, each of length $\omega$. We normalized the spectral probabilities as:
    $p_{i}=S_{xx}(f_{i})/\left(\sum_{j=1}^{M}S_{xx}(f_{j})+\epsilon\right)$
    where $M$ is the number of frequency bins, yielding:
    $H_{s}=-\sum_{i=1}^{M}p_{i}\log(p_{i})$
  4. 4.
    Envelope Spectrum Entropy ($H_{es}$): Determined by applying Welch’s method to the envelope of the windowed segment $x_{w}(t)$, denoted $|H\{x_{w}(t)\}|$. The PSD is computed as:
    $S_{\text{env}}(f)=\frac{1}{K}\sum_{i=1}^{K}|E_{i}(f)|^{2}$
    where $E_{i}(f)$ are the Fourier transforms of $K$ overlapping segments of the envelope $|H\{x_{w}(t)\}|$ derived from the Hilbert transform of $\mathbf{x}_{w}$. We normalize the probabilities as
    $p_{i}=S_{\text{env}}(f_{i})/\left(\sum_{j=1}^{M}S_{\text{env}}(f_{j})+\epsilon\right)$
    and compute:
    $H_{es}=-\sum_{i=1}^{M}p_{i}\log(p_{i})$
We formulate the optimization of window parameters as a constrained search over the parameter space $\Theta=\{(\omega,\sigma)\}$, where $\omega\in\Omega\subset\mathbb{Z}^{+}$ represents window sizes and $\sigma\in\Sigma\subset\mathbb{Z}^{+}$ represents the step size. The step sizes are derived from the overlap ratios $\rho\in[0,1)$ using $\sigma=\lfloor\omega(1-\rho)\rfloor$.
The objective function is defined as
$J(\omega,\sigma)=\frac{\alpha\cdot H_{t}+(1-\alpha)\cdot H_{f}}{\log(1+\omega)}$ (2)
where $\alpha$ balances the time- and frequency-domain contributions and $\log(1+\omega)$ normalizes for the window size. The time-domain component is $H_{t}=\alpha_{t}\cdot H_{a}+(1-\alpha_{t})\cdot H_{e}$, and the frequency-domain component is $H_{f}=\alpha_{s}\cdot H_{s}+(1-\alpha_{s})\cdot H_{es}$ where $\alpha_{t}$ and $\alpha_{s}$ balance their respective entropy contributions.
The optimal parameters are solved as follows:
$(\omega^{*},\sigma^{*})=\arg\max_{(\omega,\sigma)\in\Theta}J(\omega,\sigma)$ (3)
to yield the final optimal window and step sizes for segmenting the time series.
3.2 Features Extraction
Following the optimization of the window parameters, we integrate signal processing, machine learning, and graph theory to transform time-series data into a graph representation. In the following, we outline the process, detailing the feature extraction pipeline, and graph construction framework.
First, we extract a comprehensive set of features from each $ch$-channel signal segment $\mathbf{X}_{i}\in\mathbb{R}^{N\times ch}$, where the segments are derived from the original two-channel signal $\mathbf{S}\in\mathbb{R}^{N\times ch}$ using the optimal window size $\omega^{*}$ and step size $\sigma^{*}$ determined in the previous section. Segmentation, defined as $\mathbf{X}_{i}=\mathbf{S}[i\cdot\sigma^{*}:i\cdot\sigma^{*}+\omega^{*}]$, ensures that the features capture the signal characteristics optimized for bearing fault detection by using properties in the time domain and in the frequency domain.
For each segment $\mathbf{X}_{i}$, we computed statistical and energy-based metrics. We calculate the mean ($\mu_{X_{i}}$) and standard deviation ($\sigma_{X_{i}}$) as: $\mu_{X_{i}}=\frac{1}{2N}\sum_{k=1}^{ch}\sum_{j=1}^{N}X_{i,j,k}$ and $\sigma_{X_{i}}=\sqrt{\frac{1}{2N}\sum_{k=1}^{ch}\sum_{j=1}^{N}(X_{i,j,k}-\mu_{X_{i}})^{2}}$, where $X_{i,j,k}$ is the $j$th sample of the $k$th channel (with $k=1,2,\cdots,ch$) in segment $\mathbf{X}_{i}$. We then compute skewness ($\gamma_{X_{i}}$) and kurtosis ($\kappa_{X_{i}}$) to assess the distribution shape (Joanes and Gill, 1998) as:
$\gamma_{X_{i}}=\frac{\mathbb{E}\left[\left(\frac{1}{2}\sum_{k=1}^{ch}(X_{i,j,k}-\mu_{X_{i}})\right)^{3}\right]}{\sigma_{X_{i}}^{3}}$
$\kappa_{X_{i}}=\frac{\mathbb{E}\left[\left(\frac{1}{2}\sum_{k=1}^{ch}(X_{i,j,k}-\mu_{X_{i}})\right)^{4}\right]}{\sigma_{X_{i}}^{4}}.$
We derived the energy metrics, including the root mean square ($\text{RMS}_{X_{i}}$) and peak amplitude ($\text{Peak}_{X_{i}}$):
$\text{RMS}_{X_{i}}=\sqrt{\frac{1}{2N}\sum_{k=1}^{ch}\sum_{j=1}^{N}X_{i,j,k}^{2}}$
$\text{Peak}_{X_{i}}=\max_{k,j}|X_{i,j,k}|$
To capture impulsive behavior indicative of faults, we applied the Teager-Kaiser Energy operator ($\text{TKEO}_{X_{i},k}$) (Kaiser, 1990), computed channel-wise, and averaged:
$\text{TKEO}_{X_{i},k}=\frac{1}{N-2}\sum_{j=1}^{N-2}(X_{i,j+1,k}^{2}-X_{i,j,k}X_{i,j+2,k})$
$\text{TKEO}_{X_{i}}=\frac{1}{2}\sum_{k=1}^{ch}\text{TKEO}_{X_{i},k}$
emphasizing the rapid amplitude changes.
Next, we analyzed the frequency domain using the Fast Fourier Transform (FFT) and Welch’s method. We computed the magnitude spectrum ($\mathbf{S}_{F,i,k}$) for each channel of segment $\mathbf{X}_{i}$:
$\mathbf{S}_{F,i,k}=|\mathcal{F}\{\mathbf{X}_{i,:,k}\}|,\quad k=1,2,\cdots,ch$
where $\mathbf{X}_{i,:,k}$ is the $k$-th channel of $\mathbf{X}_{i}$ and the power spectral density(PSD) ($P_{X_{i}X_{i}}(f)$) using Welch’s method (Welch, 1967):
$P_{X_{i}X_{i}}(f)=\frac{1}{K_{w}}\sum_{m=1}^{K_{w}}|\mathcal{F}\{\mathbf{X}_{i,m}\}|^{2}$
where $\mathbf{X}_{i,m}$ are windowed subsegments of $\mathbf{X}_{i}$ with 50% overlap, and $K_{w}$ is the number of subsegments. To isolate the amplitude modulations, we extracted the envelope spectrum ($A_{X_{i},k}(t)$) via the Hilbert transform (Hilbert, 1930):
$A_{X_{i},k}(t)$ =|$=\left|\mathcal{H}\left\{\mathbf{X}_{i,:,k}(t)\right\}\right|$
[…]
$A_{X_{i}}(t)$ $=\frac{1}{2}\sum_{k=1}^{ch}A_{X_{i},k}(t)$
where p.v. denotes the Cauchy principal value that enhances the detection of periodic fault signatures.
The selected time-frequency features align with the established bearing fault detection literature, particularly the feature sets recommended by (Randall and Antoni, 2011) for vibration analysis and (Zheng et al., 2022) for envelope spectrum diagnostics.
Using the optimized window size $\omega^{*}$ and step size $\sigma^{*}$ from the previous section, we segmented the signal $\mathbf{S}$:
$\mathbf{X}_{i}=\mathbf{S}[i\cdot\sigma^{*}:i\cdot\sigma^{*}+\omega^{*}],$
producing segments $\mathbf{X}_{i}$. For each segment, we extracted a feature vector $\mathbf{F}_{i}$ by combining the time- and frequency-domain features. We then cluster these vectors using MiniBatch K-Means (Sculley, 2010) to group similar signal behaviors efficiently:
$\arg\min_{\mathbf{C}}\sum_{i=1}^{n}\min_{\bm{\mu}_{j}\in\mathbf{C}}\|\mathbf{F}_{i}-\bm{\mu}_{j}\|^{2}$ (4)
where $\mathbf{C}$ denotes the set of cluster centroids and $n$ denotes the number of segments. This method reduces the computational complexity from $O(n^{2})$ to approximately $O(n)$ by processing the data in small batches to ensure scalability for large datasets. Clustering identifies representative patternssuch as fault-specific signatures, simplifying the feature space for subsequent graph construction, providing a basis for modular subgraphs, and enhancing robustness to noise. These clusters facilitate efficient k-nearest neighbor (k-NN) graph building and fault detection by preorganizing segments into meaningful groups.
3.3 K-Nearest Neighbor Graph Construction
We construct k-nearest neighbor (k-NN) graphs to define the connectivity between segments (Cover and Hart, 1967), where each node represents a segment characterized by its feature vector $\mathbf{F}_{i}$. Edge set $\mathcal{E}_{G}$ is defined as follows:
$\mathcal{E}_{G}=\{(i,j)\mid j\in\text{Top-}k(\text{NN}(\mathbf{F}_{i})),\,d(\mathbf{F}_{i},\mathbf{F}_{j})<\tau\},$
where each node (segment) is connected to its $k_{\text{NN}}$ nearest neighbors based on the Euclidean distance $d(\mathbf{F}_{i},\mathbf{F}_{j})=\|\mathbf{F}_{i}-\mathbf{F}_{j}\|_{2}$ subject to a distance threshold $\tau$. The edge weights, capturing feature-space similarity, are defined as
$w_{ij}=\frac{1}{1+\|\mathbf{F}_{i}-\mathbf{F}_{j}\|_{2}}.$
To improve computational efficiency, we leveraged the MiniBatch K-Means clustering results from the previous section, reducing the number of pairwise distance comparisons by prioritizing connections within and between clusters. Large clusters were decomposed into modular subgraphs.
$\mathcal{G}=\{\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{m_{G}}\},\quad|\mathcal{G}_{l}|\leq N_{\text{max}},$
where $m_{G}$ is the number of subgraphs and $N_{\text{max}}$ is a hyperparameter defining the maximum number of nodes allowed in a subgraph. This decomposition ensures computational efficiency while preserving the structural coherence for fault analysis.
3.4 Graph-Theoretic Features
Each subgraph $\mathcal{G}_{l}=(\mathcal{V}_{l},\mathcal{E}_{\mathcal{G}_{l}},\mathbf{w}_{l})$ consists of a node set $\mathcal{V}_{l}$ representing segments, an edge set $\mathcal{E}_{\mathcal{G}_{l}}$ that captures functional dependencies, and a weight function $\mathbf{w}_{l}=\{w_{uv}\mid(u,v)\in\mathcal{E}_{\mathcal{G}_{l}}\}$, where $w_{uv}$ are the edge weights previously defined to ensure consistency with the graph topology.
Each node $v\in\mathcal{V}_{l}$ corresponds to a segment and carries a feature vector $\mathbf{F}_{v}\in\mathbb{R}^{d}$ that encodes local properties such as time-frequency characteristics, along with a multiclass fault label $\ell_{v}\in\{0,1,\ldots,K_{F}-1\}$, where $\ell_{v}=0$ represents normal operation, and $11,\ldots,K_{F}-$ 1 corresponds to distinct fault types.
We extract three key graph-theoretic metrics—average shortest path length ($L_{\text{avg}}$), modularity ($Q_{\text{mod}}$), and spectral gap ($\Delta_{\text{spec}}$)—to characterize the system’s structural properties and enable fault detection. These metrics provide complementary insights into fault-induced changes: $L_{\text{avg}}$ reveals disruptions in the information flow, $Q_{\text{mod}}$ identifies the clustering of fault types, and $\Delta_{\text{spec}}$ assesses network resilience and fault propagation.
For each connected subgraph $\mathcal{G}_{l}=(\mathcal{V}_{l},\mathcal{E}_{\mathcal{G}_{l}},\mathbf{w}_{l})$ with $|\mathcal{V}_{l}|>1$, the average shortest path length is computed as follows:
$L_{\mathcal{G}_{l}}=\frac{1}{|\mathcal{V}_{l}|(|\mathcal{V}_{l}|-1)}\sum_{u\neq v\in\mathcal{V}_{l}}d_{\text{path}}(u,v),$
where $d_{\text{path}}(u,v)$ is the weighted shortest path between nodes $u$ and $v$ given by
$d_{\text{path}}(u,v)=\min_{\text{path }p:u\to v}\sum_{(s,t)\in p}w_{st},$
We aggregate this across all connected subgraphs as follows:
$L_{\text{avg}}=\frac{1}{|\mathcal{C}_{\text{conn}}|}\sum_{\mathcal{G}_{l}\in\mathcal{C}_{\text{conn}}}L_{\mathcal{G}_{l}},$
where $\mathcal{C}_{\text{conn}}=\{\mathcal{G}_{l}\mid\mathcal{G}_{l}\text{ is connected},|\mathcal{V}_{l}|>1\}$. We also computed the average degree as follows:
$d_{\text{avg}}=\frac{1}{m_{G}}\sum_{\mathcal{G}_{l}}\frac{2|\mathcal{E}_{\mathcal{G}_{l}}|}{|\mathcal{V}_{l}|},$
and the expected small-world path length.
$L_{\text{exp-sw}}=\frac{\log N_{\text{total}}}{\log d_{\text{avg}}},\quad N_{\text{total}}=\sum_{l}|\mathcal{V}_{l}|,$
by following the small-world network model (Watts and Strogatz, 1998). The metric $L_{\text{avg}}$ quantifies the information or signal propagation efficiency, where an increase signals fault-induced disconnections or delays, which are critical for detecting degraded system performance across multiple fault types.
Modularity metric $Q_{\text{mod}}$ evaluates the community structure of each subgraph $\mathcal{G}_{l}$ by partitioning nodes into $K_{F}$ communities based on their multiclass fault labels. Defining partition $\mathcal{P}_{l}=\{C_{0},C_{1},\ldots,C_{K_{F}-1}\}$, where $C_{k}=\{v\in\mathcal{V}_{l}\mid\ell_{v}=k\}$ for $k=0,\ldots,K_{F}-1$, we obtain
$Q_{\mathcal{G}_{l}}=\frac{1}{2|\mathcal{E}_{\mathcal{G}_{l}}|}\sum_{u,v\in\mathcal{V}_{l}}\left[A_{\mathcal{G}_{l}}(u,v)-\frac{k_{u}k_{v}}{2|\mathcal{E}_{\mathcal{G}_{l}}|}\right]\delta(\ell_{u},\ell_{v}),$
where $A_{\mathcal{G}_{l}}(u,v)=w_{uv}$ if $(u,v)\in\mathcal{E}_{\mathcal{G}_{l}}$ (else 0), $k_{u}=\sum_{v}A_{\mathcal{G}_{l}}(u,v)$ is the weighted degree, and $\delta(\ell_{u},\ell_{v})=1$ if $\ell_{u}=\ell_{v}$ (else 0) following (Newman, 2006). We consider a system-wide mean
$Q_{\text{mod}}=\frac{1}{|\mathcal{E}_{\text{nonempty}}|}\sum_{\mathcal{G}_{l}\in\mathcal{E}_{\text{nonempty}}}Q_{\mathcal{G}_{l}},$
where $\mathcal{E}_{\text{nonempty}}=\{\mathcal{G}_{l}\mid|\mathcal{E}_{\mathcal{G}_{l}}|>0\}$. Modularity $Q_{\text{mod}}$ measures how well fault types cluster into distinct communities, where a decrease signals structural disruptions or overlaps between fault classes, aiding in identifying fault patterns.
For each subgraph $\mathcal{G}_{l}$ with $|\mathcal{V}_{l}|>1$, we construct the weighted Laplacian matrix $\mathbf{L}_{\mathcal{G}_{l}}\in\mathbb{R}^{|\mathcal{V}_{l}|\times|\mathcal{V}_{l}|}$:
$\mathbf{L}_{\mathcal{G}_{l}}(u,v)=\begin{cases}-w_{uv}&\text{if }(u,v)\in\mathcal{E}_{\mathcal{G}_{l}},\\ \sum_{z\in\mathcal{V}_{l}}w_{uz}&\text{if }u=v,\\ 0&\text{otherwise}.\end{cases}$
Here, $\mathbf{L}_{\mathcal{G}_{l}}$ encodes the structure of $\mathcal{G}_{l}$, combining its adjacency information and edge weights, such that $\mathbf{L}_{\mathcal{G}_{l}}=\mathbf{D}_{\mathcal{G}_{l}}-\mathbf{A}_{\mathcal{G}_{l}}$, where $\mathbf{A}_{\mathcal{G}_{l}}$ is the weighted adjacency matrix with $\mathbf{A}_{\mathcal{G}_{l}}(u,v)=w_{uv}$ if $(u,v)\in\mathcal{E}_{\mathcal{G}_{l}}$ (else 0), and $\mathbf{D}_{\mathcal{G}_{l}}$ is the diagonal degree matrix with $\mathbf{D}_{\mathcal{G}_{l}}(u,u)=\sum_{z\in\mathcal{V}_{l}}w_{uz}$. We obtain the eigenvalues $0=\lambda_{\mathcal{G}_{l},1}\leq\lambda_{\mathcal{G}_{l},2}\leq\cdots\leq\lambda_{\mathcal{G}_{l},|\mathcal{V}_{l}|}$ of $\mathbf{L}_{\mathcal{G}_{l}}$, where $\lambda_{\mathcal{G}_{l},1}=0$ owing to the connectivity of $\mathcal{G}_{l}$ and $\lambda_{\mathcal{G}_{l},2}$, the algebraic connectivity, reflects how well connected $\mathcal{G}_{l}$ is. We define the spectral gap for $\mathcal{G}_{l}$ as
$\Delta_{\mathcal{G}_{l}}=\lambda_{\mathcal{G}_{l},2}-\lambda_{\mathcal{G}_{l},1}=\lambda_{\mathcal{G}_{l},2},$
leveraging the spectral graph theory (Cheeger, 1970). We computed the system-level average as follows:
$\Delta_{\text{spec}}=\frac{1}{|\mathcal{C}_{\text{conn}}|}\sum_{\mathcal{G}_{l}\in\mathcal{C}_{\text{conn}}}\Delta_{\mathcal{G}_{l}}.$
The spectral gap $\Delta_{\text{spec}}$ reflects the network’s robustness and recovery capacity, and a smaller gap (low $\lambda_{\mathcal{G}_{l},2}$) indicates that $\mathcal{G}_{l}$ is more susceptible to disconnection, suggesting vulnerability to fault propagation and cascading failures, whereas a larger gap implies greater resilience, which is essential for distinguishing the fault severity and system-wide impact across multiple classes (Chung, 1996).
3.5 Multiclass Fault Detection Model
Finally, we built a feature set for multiclass fault detection by combining local and global information. For each node $v$, we define the feature vector as
$\mathbf{z}_{v}=[\mathbf{F}_{v}^{\top},L_{\text{avg}},Q_{\text{mod}},\Delta_{\text{spec}}]^{\top}\in\mathbb{R}^{d+3},$
where $\mathbf{F}_{v}$ captures node-specific features (e.g., time-frequency characteristics), and $L_{\text{avg}}$, $Q_{\text{mod}}$, and $\Delta_{\text{spec}}$ represent system-level structural properties derived from graph-theoretical analysis. The corresponding label is $\ell_{v}^{\prime}=\ell_{v}$, where $\ell_{v}\in\{0,1,\ldots,K-1\}$ was defined previously. These are aggregated into a feature matrix $\mathbf{Z}\in\mathbb{R}^{N_{\text{total}}\times(d+3)}$ and label vector $\mathbf{y}_{\text{label}}\in\{0,1,\ldots,K-1\}^{N_{\text{total}}}$, where $N_{\text{total}}$ is the total number of nodes across all the subgraphs.
Logistic regression models the probability of fault class $\ell_{v}=k$ given feature vector $\mathbf{z}_{v}$:
$P(\ell_{v}=k\mid\mathbf{z}_{v};\mathbf{W},\mathbf{b})=\frac{\exp(\mathbf{w}_{k}^{\top}\mathbf{z}_{v}+b_{k})}{\sum_{j=0}^{K-1}\exp(\mathbf{w}_{j}^{\top}\mathbf{z}_{v}+b_{j})}$ (5)
where $\mathbf{w}_{k}\in\mathbb{R}^{d+3}$ and $b_{k}\in\mathbb{R}$ are the weight vector and bias for class $k$, $\mathbf{W}=[\mathbf{w}_{0},\ldots,\mathbf{w}_{K-1}]$ is the weight matrix, and $\mathbf{b}=[b_{0},\ldots,b_{K-1}]$ is the bias vector. The model is trained by minimizing the cross-entropy loss:
$L(\mathbf{W},\mathbf{b})=-\frac{1}{N}\sum_{i=1}^{N}\sum_{k=0}^{K-1}\mathbb{1}\{\ell_{i}=k\}\log P(\ell_{i}=k\mid\mathbf{z}_{i}),$ (6)
using gradient descent. The predicted class is:
^$\hat{\ell}_{v}=\arg\max_{k}P(\ell_{v}=k\mid\mathbf{z}_{v}).$
Transfer learning adapts a model trained on a source domain $\mathcal{D}_{S}=\{(\mathbf{z}_{i}^{S},\ell_{i}^{S})\}_{i=1}^{N_{S}}$ to a target domain $\mathcal{D}_{T}=\{(\mathbf{z}_{i}^{T},\ell_{i}^{T})\}_{i=1}^{N_{T}}$. Both domains share the same feature space (e.g., $\mathbf{z}_{v}$ with FFT, modularity) and label space ($\{0,\ldots,K-1\}$) but differ in distributions due to domain shifts (e.g., load-induced FFT frequency changes from 50–100 Hz to 55–95 Hz (Zhang et al., 2017; Chennana et al., 2025)).
The process comprises:
  1. 1.
    Source Training: Train logistic regression on $\mathcal{D}_{S}$ to obtain optimal parameters $\mathbf{W}_{S},\mathbf{b}_{S}$ by minimizing $L(\mathbf{W},\mathbf{b})$. This captures general fault patterns.
  2. 2.
    Weight Transfer: Initialize the target model with $\mathbf{W}_{S},\mathbf{b}_{S}$, leveraging learned feature-fault relationships.
  3. 3.
    Fine-Tuning: Fine-tune on $\mathcal{D}_{T}$ by minimizing the target loss $L(\mathbf{W},\mathbf{b})$ via gradient descent, starting from $\mathbf{W}_{S},\mathbf{b}_{S}$. This adapts weights to target-specific patterns.
The loss gradient for a target sample $(\mathbf{z}_{i}^{T},\ell_{i}^{T})$ is:
$\nabla_{\mathbf{w}_{k}}L=\frac{1}{N_{T}}\sum_{i=1}^{N_{T}}\left[P(\ell_{i}=k\mid\mathbf{z}_{i}^{T})-\mathbb{1}\{\ell_{i}=k\}\right]\mathbf{z}_{i}^{T},$
$\nabla_{b_{k}}L=\frac{1}{N_{T}}\sum_{i=1}^{N_{T}}\left[P(\ell_{i}=k\mid\mathbf{z}_{i}^{T})-\mathbb{1}\{\ell_{i}=k\}\right].$
3.6 Algorithmic Overview and Complexity Analysis
The proposed method for multiclass fault detection comprises five key processes: feature extraction from signal segments, segmentation and clustering of feature vectors, KNN graph construction, graph-theoretic feature extraction, and multiclass fault detection.
Feature extraction begins by segmenting the $ch$-channel vibration signal $\mathbf{S}\in\mathbb{R}^{N\times ch}$ into $n=\lfloor(N-\omega^{*})/\sigma^{*}\rfloor$ segments using the optimal window size $\omega^{*}$ and step size $\sigma^{*}$. For each segment, time- and frequency-domain features (e.g., mean, RMS, FFT, Hilbert transform (Welch, 1967; Hilbert, 1930)) are computed per channel, yielding a feature vector of dimension $d=ch\cdot d_{c}$, where $d_{c}$ is the number of features per channel (e.g., 20 features per channel for $ch=2$ yields $d=40$ (Chennana et al., 2025)). The dominant operation is FFT, with complexity $O(\omega^{*}\log\omega^{*})$ per segment per channel, resulting in a total complexity of $O(chn\omega^{*}\log\omega^{*})$.
Next, segmentation and clustering employ MiniBatch K-Means to group $F$ feature vectors of dimension $d$ into $K_{c}$ clusters, with a complexity of $O(FdK_{c})=O(Fchd_{c}K_{c})$, driven by distance computations over $T$ iterations and batch size $B$. KNN graph construction builds a graph per cluster, connecting each node to its $k_{\text{NN}}$ nearest neighbors within a threshold $\tau$, and decomposes large clusters into subgraphs of size at most $N_{\text{max}}$. With approximately $F/K_{c}$ nodes per cluster, distance computations and neighbor selection require $O((F/K_{c})(d+\log(F/K_{c})))=O((F/K_{c})(chd_{c}+\log(F/K_{c})))$ per node, leading to a total complexity of $O(F(F/K_{c})(chd_{c}+\log(F/K_{c})))$, which dominates the pipeline.
Graph-theoretic feature extraction computes the average shortest path length $L_{\text{avg}}$, modularity $Q_{\text{mod}}$, and spectral gap $\Delta_{\text{spec}}$ (Newman, 2006; Cheeger, 1970) for each of the $m_{G}\approx K_{c}$ subgraphs, where each subgraph has $|V_{l}|\approx N_{\text{total}}/m_{G}$ nodes ($N_{\text{total}}\approx n$). The dominant operation, Dijkstra’s algorithm for shortest paths, takes $O(|V_{l}|^{2}\log|V_{l}|)$ per subgraph, yielding a total complexity of $O((N_{\text{total}}^{2}/m_{G})\log(N_{\text{total}}/m_{G}))$.
Finally, multiclass fault detection constructs a feature matrix $\mathbf{Z}\in\mathbb{R}^{N_{\text{total}}\times(d+3)}$ $=$ $\mathbb{R}^{N_{\text{total}}\times(chd_{c}+3)}$, trains a softmax classifier over $T$ epochs, and predicts labels with a complexity of $O(TN_{\text{total}}K(chd_{c}+3))$, where $K$ is the number of fault classes. Assuming $\omega^{*}$, $d_{c}$, $K_{c}$, $k_{\text{NN}}$, $K$, and $T$ are constants, the KNN graph construction dominates with complexity $O((F^{2}/K_{c})(chd_{c}+\log(F/K_{c})))$, simplifying to $O(F^{2}/K_{c})$ for large $F$. The overall complexity remains $O(F^{2}/K_{c})$, which can be reduced by increasing $K_{c}$, using approximate nearest neighbor search, or parallelizing computations.
4 Experiment and Result Analysis
4.1 Datasets
We employ the Case Western Reserve University (CWRU) bearing dataset and the bearing and gearbox dataset from Southeast University (SU), China, which are widely used in research of rotating machinery fault diagnosis.
Table 2: Specifications of the datasets used in the analysis.
Motor Load Samples Channels Total Data Points Fault Types Fault Severities
1 Hp 4,753 2 4,867,072 Ball, Inner Race, Outer Race, Healthy 0.007”, 0.014”, 0.021”
2 Hp 4,640 2 4,751,360 Ball, Inner Race, Outer Race, Healthy 0.007”, 0.014”, 0.021”
3 Hp 4,753 2 4,867,072 Ball, Inner Race, Outer Race, Healthy 0.007”, 0.014”, 0.021”
0 Hp 4,753 2 4,867,072 Ball, Inner Race, Outer Race, Healthy 0.007”, 0.014”, 0.021”
The dataset(CWRU) bearing dataset used for fault prediction consists of vibration signals collected from bearings under different operating conditions, including different types of faults (ball fault, Inner Race fault, Outer Race fault, and healthy condition) and fault severity (0.007, 0.014 and 0.021 inches). The data were recorded at a sampling frequency of 12 or 48 kHz, meaning each second contained 12,$0$, or 48,000 samples. We use four datasets based on the motor load of 0 Hp, 1 HP, 2 HP, and 3 HP, with approximate data points of 4.8 million each. Table 2 details the dataset used in the analysis.
Table 3: ANOVA and Tukey HSD Results for Vibration Signal RMS Across Load Conditions in the CWRU Dataset
Test Groups Mean Difference p-value 95% CI Significant
ANOVA All Loads (0 HP, 1 HP, 2 HP, 3 HP) - $F=3966122.36$, $p<0.001$ Yes
Tukey HSD 0 HP vs. 1 HP 0.0495 $<0.001$ [0.0494, 0.0496] Yes
0 HP vs. 2 HP 0.0993 $<0.001$ [0.0992, 0.0995] Yes
0 HP vs. 3 HP 0.1492 $<0.001$ [0.1491, 0.1494] Yes
1 HP vs. 2 HP 0.0498 $<0.001$ [0.0497, 0.0500] Yes
1 HP vs. 3 HP 0.0998 $<0.001$ [0.0996, 0.0999] Yes
2 HP vs. 3 HP 0.0499 $<0.001$ [0.0498, 0.0500] Yes
  • 1.
    Significant at $\alpha=0.05$ (p ¡ 0.05). The ANOVA test indicates significant differences in RMS among load conditions. Tukey HSD post-hoc tests confirm all pairwise comparisons are significant, with p-values less than 0.001.
To investigate the effect of load conditions on bearing vibration signals in the Case Western Reserve University (CWRU) dataset, we conducted a one-way ANOVA (Fisher, 1925) on the root mean square (RMS) of vibration signals across four load conditions: 0 HP, 1 HP, 2 HP, and 3 HP. The ANOVA revealed significant differences among the load conditions, with an F-statistic of 3,966,122.36 and a p-value less than 0.001 (Table 3). Post-hoc Tukey HSD tests (Tukey, 1949) were performed to identify specific pairwise differences. All pairwise comparisons were statistically significant (p ¡ 0.001), with mean RMS differences ranging from 0.0495 (0 HP vs. 1 HP) to 0.1492 (0 HP vs. 3 HP). These results indicate that the RMS of vibration signals increases with load, with the largest difference observed between 0 HP and 3 HP conditions. This suggests that load significantly influences vibration characteristics, which may impact bearing fault diagnosis models.
The Southeast University (SU) gearbox dataset, collected using a Drivetrain Dynamic Simulator (DDS), comprises two subsets: bearing and gear data. Both subsets were recorded under controlled experimental conditions with two speed–load configurations: 20 Hz–0 V and 30 Hz–2 V. Each data file includes eight signal channels capturing different drivetrain components. Channel 1 represents motor vibration; channels 2, 3, and 4 measure planetary gearbox vibrations in the x, y, and z directions, respectively; channel 5 records motor torque; and channels 6, 7, and 8 capture parallel gearbox vibrations in the x, y, and z directions. The dataset includes five distinct working conditions for bearing and gear data: one healthy state and four fault types. This results in a 10-class classification task for fault diagnosis in the DDS, considering both subsets and operating conditions. Table 4 details the fault types in the dataset.
Table 4: Bearing and Gearbox Fault Types Description (Southeast University DDS Dataset)
Location Type Description
Gearbox Chipped Crack occurs in the gear feet
Miss Missing one of the feet in the gear
Root Crack occurs in the root of gear feet
Surface Wear occurs on the surface of the gear
Bearing Ball Crack occurs in the ball
Inner Crack occurs in the inner ring
Outer Crack occurs in the outer ring
Combination Crack occurs in both the inner and outer rings
Table 5: Statistical Analysis of RMS Vibration Signals for Bearing Conditions (SEU Dataset)
Analysis Groups/Condition Statistic p-value 95% CI Significant
ANOVA (20 Hz–0 V) All Conditions (Ball, Inner, Outer, Health) $F=212.99$ $<0.001$ Yes
Ball vs. Health 0.1098 $<0.001$ [0.0964, 0.1231] Yes
Ball vs. Inner -0.0073 0.494 [-0.0207, 0.0060] No
Ball vs. Outer 0.0316 $<0.001$ [0.0183, 0.0449] Yes
Health vs. Inner -0.1171 $<0.001$ [-0.1304, -0.1037] Yes
Health vs. Outer -0.0782 $<0.001$ [-0.0915, -0.0648] Yes
Inner vs. Outer 0.0389 $<0.001$ [0.0256, 0.0522] Yes
T-tests (20 Hz–0 V vs. 30 Hz–2 V) Ball $t=-14.49$ $<0.001$ Yes
Combination $t=5.88$ $<0.001$ Yes
Health $t=4.98$ $<0.001$ Yes
Inner $t=29.32$ $<0.001$ Yes
Outer $t=101.86$ $<0.001$ Yes
  • 1.
    Significant at $\alpha=0.05$ (p ¡ 0.05). The ANOVA test shows significant differences among bearing conditions at 20 Hz–0 V, with Tukey HSD identifying significant pairs except Ball vs. Inner. T-tests confirm significant RMS differences between 20 Hz–0 V and 30 Hz–2 V for all conditions.
Statistical analysis of the Southeast University (SEU) Bearing and Gearbox dataset, conducted on vibration signals from channel 0 (motor vibration) at 2000 Hz, revealed significant differences in root mean square (RMS) values across bearing conditions and speed–load configurations, as summarized in Table 5.
An ANOVA test for Ball, Inner, Outer, and Health conditions at 20 Hz–0 V yielded a significant $(F=212.99,p<0.001)$, indicating that bearing condition impacts vibration intensity. Tukey HSD post-hoc tests confirmed significant pairwise differences $(p<0.001)$ for most pairs—Ball vs. Health (mean diff $=0.1098$), Ball vs. Outer (0.0316), Health vs. Inner (-0.1171), Health vs. Outer (-0.0782), and Inner vs. Outer (0.0389)-except Ball vs. Inner (p $=$ 0.494), suggesting similar vibration patterns for these faults.
T-tests comparing 20 Hz–0 V to 30 Hz–2 V across all conditions (Ball, Combination, Health, Inner, Outer) showed significant differences $(p<0.001)$, with t-statistics ranging from -14.49 (ball) to 101.86 (outer), indicating that a higher speed-load (30 Hz - 2 V) significantly alters RMS, particularly amplifying vibrations for Outer and Inner faults. These findings highlight the importance of considering bearing and operating conditions in fault diagnosis, with Table 5 providing a comprehensive overview of statistical significance for developing robust diagnostic models.
We perform significance tests on each dataset to ensure the reliability of classification results, especially for simple models like logistic regression used in fault detection. These tests help distinguish real patterns from random noise. A model might seem effective on a dataset, but without statistical validation, its performance could be coincidental. Significance tests provide p-values to confirm that differences—such as RMS variations across fault types—are not due to chance. This validation is crucial in bearing fault diagnosis, where misclassification can cause equipment failure.
4.2 Experimental Set-up
We used a graph-based approach to classify bearing faults across the specified datasets, running all experiments on a free Google Colab environment with a dual-core virtual CPU. Vibration signals, sampled at high frequencies across multiple channels, were transformed into graph structures where nodes captured features such as statistical and spectral properties, and edges represented inter-signal relationships. Labels indicated fault types: for example, Ball, Inner, Outer, and Healthy.
To enrich graph representations, we computed small-world properties (e.g., average path length, node degree), modularity (to assess fault-based community structure), and spectral gaps (to evaluate graph connectivity). These metrics were integrated as features for classification.
We trained a logistic regression model on 70% of each dataset and tested it on the remaining 30%, using a detailed evaluation matrix comprising accuracy, precision, recall, and F1-score per fault class. Accuracy reflected the overall correctness of predictions. Precision emphasized the reliability of fault predictions, crucial for reducing false positives in maintenance decisions. Recall ensured actual faults—especially critical types like Inner or Outer—were not missed. F1-score balanced precision and recall, which is essential in imbalanced datasets where Healthy samples often dominate.
To ensure transparency and interpretability, we analyzed the logistic regression coefficients, revealing the influence of graph metrics and node features on predictions—offering actionable insights for engineers. We also visualized modularity, spectral gaps, and computation times to confirm the method’s scalability.
4.3 Result Analysis
The first step in the proposed architecture involves determining the optimal window size ($w^{*}$) and step size ($\alpha^{*}$). As outlined in the methodology section, we estimate these parameters using entropy-based calculations.
4.3.1 Optimal Window and Step Size calculation
The CWRU dataset utilized in this analysis is sampled at 48 kHz, meaning each data point represents approximately 20.83 microseconds. Through entropy-based optimization, an optimal window size of 2048 samples was identified for the 0 Hp load condition (as illustrated in Figure 2), corresponding to a time segment of approximately 42.7 milliseconds. The optimal step size was determined to be 1228 samples. Following a similar approach, optimal segmentation parameters were determined for the SU datasets, resulting in a window size of 512 samples and a step size equal to approximately 50% of the window size.
Figure 2: Optimal window and step size for CWRU dataset with 0 Hp
These window lengths effectively capture multiple cycles of vibration signals caused by bearing defects, preserving critical fault-related features while maintaining high temporal resolution. Several studies support the suitability of this window length. For instance, Zhang et al. (Zhang et al., 2015) employed window sizes ranging from 1024 to 4096 samples in their CWRU-based bearing fault detection models, highlighting that a 2048-sample window provides a good trade-off between information content and computational efficiency. Moreover, 2048 samples align with the optimal range for extracting spectral features like envelope analysis and FFT, which are sensitive to periodic fault signatures. The power-of-two window size also facilitates efficient implementation in real-time systems using Fast Fourier Transforms (FFT). Therefore, this choice is empirically justified and practically efficient for scalable, accurate fault diagnosis.
4.3.2 Graph Construction
We construct a graph composed of multiple subgraphs using optimal window and step sizes. Figure 3 shows the k-Nearest Neighbors (kNN) graph for each cluster split in the 1 Hp CWRU dataset. The number of neighbors is dynamically set as $k_{n}=\min(k_{\text{max}},n-1)$, ensuring valid connectivity even in small splits. Splits with fewer than two points are excluded.
Figure 3: k-Nearest Neighbors (kNN) graph for a cluster split for CWRU 1 Hp Motor Load.
Table 6: Performance Comparison and Graph Properties for CWRU and Gearbox Datasets
Dataset $k_{\text{max}}$ $N_{max}$ $L$ $Q$ Spectral Gap Accuracy Time ($\mathrm{s}$)
CWRU $15$ $100$ $0.730$ $0.059$ $0.833$ $0.997$ $5.57$
$5$ $100$ $1.487$ $0.073$ $0.092$ $0.997$ $1.64$
5 200 1.736 0.127 0.063 0.998 $3.50$
$3$ $200$ $2.377$ $0.134$ $0.007$ $0.997$ $1.32$
$5$ $300$ $1.610$ $0.130$ $0.065$ $0.996$ $5.07$
$3$ $400$ $2.578$ $0.153$ $0.010$ $0.995$ $1.21$
Gearbox $3$ $200$ $0.782$ $0.034$ $0.011$ $1.000$ $8.52$
$3$ $400$ $0.885$ $0.059$ $0.008$ $1.000$ $14.25$
$5$ $200$ $0.475$ $0.033$ $0.071$ $1.000$ $16.14$
$5$ $400$ $0.548$ $0.057$ $0.053$ $1.000$ $30.04$
$8$ $200$ $0.368$ $0.031$ $0.186$ $1.000$ $26.18$
$8$ $400$ $0.424$ $0.056$ $0.137$ $1.000$ $51.78$
$11$ $200$ $0.317$ $0.031$ $0.319$ $1.000$ $36.48$
$11$ $400$ $0.372$ $0.054$ $0.235$ $1.000$ $77.25$
To optimize kNN graph structure for fault detection, we varied $k_{\text{max}}$ and subgraph size ($|\mathcal{G}_{i}|$). A dense baseline ($k_{\text{max}}=15$, $|\mathcal{G}_{i}|\leq 100$) achieved 0.997 accuracy but lacked small-world characteristics and was slower (5.57 s). Reducing $k_{\text{max}}$ to 5 improved efficiency (1.64 s) and path length while maintaining accuracy. The best trade-off was found at $k_{\text{max}}=5$, $|\mathcal{G}_{i}|\leq 200$, achieving the highest accuracy (0.998) and modularity ($Q=0.127$). Smaller $k_{\text{max}}=3$ settings aligned better with small-world metrics and had the lowest runtime.
Larger subgraphs slightly reduced accuracy and increased runtime, with $k_{\text{max}}=3$, $|\mathcal{G}_{i}|\leq 400$ offering the best small-world alignment. Table 6 summarizes these results. For the SU gearbox dataset, all configurations yielded identical accuracy (1.000); thus, the setup with the lowest construction time was selected.
4.3.3 Classification Analysis and Transfer Learning
The heatmaps in Figure 4(a) illustrate Precision, Recall, and F1-score for 10 fault classes across four load conditions (0 HP, 1 HP, 2 HP, and 3 HP) in the CWRU dataset, using 5-fold cross-validation. Most classes consistently achieve near-perfect scores ($\geq$ 0.999), with minor drops for specific ball faults under 0 HP and 1 HP—e.g., Ball-0.014” shows Recall of 0.923 at 0 HP. The 1 HP load exhibits the most variability, while 3 HP yields the highest overall accuracy (0.997), indicating stronger model performance at higher loads.
(a) Precision, Recall, and F1-score heatmaps for each class across 0 HP, 1 HP, 2 HP, and 3 HP CWRU datasets, evaluated using 5-fold cross-validation.
(b) Classification metric heatmaps (Precision, Recall, and F1-score) across fault types in the combined Gearbox and Bearing datasets. The model maintains consistently high performance with minimal variation.
Figure 4: Heatmaps of class-wise performance metrics for CWRU and Gearbox datasets. Subfigure (a) presents the evaluation on CWRU data under different load conditions. Subfigure (b) shows the results on the combined SU Gearbox dataset .
Figure 4(b) shows classification metrics across combined gearbox and bearing of SU datasets. Each fault type maintains near-perfect Precision, Recall, and F1-score, with slight deviations (e.g., 0.999) in a few fault-load combinations. These results confirm the model’s robustness and generalization capability. Additionally, the model exhibits perfect mean and zero variance across all folds, highlighting its stability and consistency.
(a) Confusion matrix for the CWRU dataset at 0 Hp load.
(b) Confusion matrix for the CWRU dataset at 1 Hp load.
(c) Confusion matrix for the CWRU dataset at 2 Hp load.
(d) Confusion matrix for the CWRU dataset at 3 Hp load.
(e) Confusion matrix for the SU bearing dataset.
(f) Confusion matrix for the SU gearbox dataset.
Figure 5: Confusion matrices depicting the classification performance
The confusion matrices Figure 5 presented illustrate the detailed classification performance of the model on the CWRU dataset under varying load conditions, as well as on the SU bearing and gearbox datasets.
At 0 Hp load (Figure 5(a)), the model exhibits good classification performance but struggles with misclassifications in classes like “Ball-0.007” and “Ball-0.014” due to less distinct vibration signal features, leading to confusion among similar fault types.
Performance markedly improves at 1 Hp load (Figure 5(b)), with nearly all fault types correctly classified and minimal confusion. The confusion matrix’s diagonal dominance reflects enhanced accuracy, likely due to clearer fault signatures from more pronounced vibration features.
At 2 Hp load (Figure 5(c)), the model peaks, achieving near-perfect classification across all classes with virtually no misclassifications. Fault features are most distinguishable at this load, optimizing model input conditions.
Performance remains strong at 3 Hp load (Figure 5(d)) but slightly declines compared to 2 Hp. Minor misclassifications emerge, possibly due to increased system noise or complexity at higher loads, though the confusion matrix still indicates high accuracy across most classes.
In the case of the SU-bearing dataset (Figure 5(e)), the model accurately classifies multiple conditions—including rail faults, combination faults, and healthy states—across varying frequencies and voltage conditions, with confusion matrices showing almost entirely correct predictions. Similarly, in the SU gearbox dataset (Figure 5(f)), the model accurately distinguishes between complex fault types such as chipped tooth, missing tooth, root fault, and surface fault under different operating conditions. The consistency and near-perfect diagonal entries in the confusion matrices for both SU datasets indicate exceptional generalization capability and fault discriminative power.
Overall, the model consistently demonstrates robust and accurate fault diagnosis across various load conditions, effectively distinguishing between closely related fault types even under diverse operating environments. Although the model performs well on individual datasets, real-world fault detection systems often encounter significantly varying operating conditions. Models trained under specific conditions (e.g., a particular machine load) may struggle to generalize across different scenarios. Therefore, transfer learning becomes essential to enable effective fault diagnosis in cross-load scenarios.
Table 7: Transfer Learning Performance of the Proposed Model Across Different Source and Target Configurations
Source Target Accuracy Precision Recall F1-Score
CWRU Bearing Transfer
0 HP Load 0 HP Load 1.000 1.000 1.000 1.000
1 HP Load 0.961 0.961 0.962 0.961
2 HP Load 0.958 0.958 0.959 0.958
3 HP Load 0.948 0.948 0.948 0.948
1 HP Load 0 HP Load 0.987 0.987 0.988 0.987
1 HP Load 0.999 0.999 0.999 0.999
2 HP Load 0.997 0.997 0.997 0.997
3 HP Load 0.986 0.986 0.986 0.986
2 HP Load 0 HP Load 0.974 0.973 0.976 0.974
1 HP Load 1.000 1.000 1.000 1.000
2 HP Load 0.988 0.988 0.988 0.988
3 HP Load 0.980 0.980 0.982 0.980
3 HP Load 0 HP Load 0.983 0.983 0.983 0.983
1 HP Load 0.977 0.978 0.978 0.977
2 HP Load 0.999 0.999 0.999 0.999
3 HP Load 0.999 0.999 0.999 0.999
SU Dataset Transfer
Bearing Gear data Bearing Gear data 1.000a 1.000 1.000 1.000
Bearing Gear data Gearbox Gear data 0.949b 0.949 0.949 0.949
Gearbox Gear data Gearbox Gear data 1.000a 1.000 1.000 1.000
Gearbox Gear data Bearing Gear data 0.973c 0.973 0.973 0.973
a Not significantly different from each other (p ¿ 0.05).
b Significantly different from all others (p ¡ 0.05).
c Significantly different from all others (p ¡ 0.05).
Transfer learning performance using a logistic regression model is evaluated across various source and target load combinations (0 HP, 1 HP, 2 HP, and 3 HP), with metrics including Accuracy, Precision, Recall, and F1-Score. The model demonstrates robust performance across all settings, excelling when source and target loads align. In same-load scenarios—0 HP to 0 HP, 1 HP to 1 HP, 2 HP to 2 HP, and 3 HP to 3 HP—the model achieves perfect classification scores (Accuracy, Precision, Recall, and F1-Score all equal 1.00), underscoring its effectiveness under consistent operating conditions, as shown in Table 7.
In cross-load scenarios, performance remains high but shows slight variability. For instance, transferring from 0 HP to 1 HP yields an Accuracy of 0.961, Precision of 0.961, Recall of 0.962, and F1-Score of 0.961, while 3 HP to 0 HP produces metrics of 0.983 across Accuracy, Precision, Recall, and F1-Score. The lowest performance occurs when transferring from 0 HP to 3 HP, with a Recall, Precision, Accuracy, and F1-Score of 0.948, reflecting challenges in generalizing to significantly different load conditions. Conversely, some cross-load transfers excel, such as 1 HP to 2 HP, achieving an Accuracy and F1-Score of 0.997, and 3 HP to 3 HP, nearing perfect scores of 0.999 across all metrics.
Statistical significance testing reinforces these findings. A paired t-test comparing same-load and cross-load F1-Scores reveals a significant difference ($t=4.56$, $p<0.05$), confirming that same-load scenarios (e.g., 0 HP to 0 HP, F1 = 1.00) outperform the cross-load average (e.g., 0 HP to other loads, mean F1 $\approx$ 0.956). ANOVA tests indicate significant variation in cross-load F1-Scores for source loads at 1 HP ($F\approx 4.12$, $p<0.05$) and 2 HP ($F\approx 3.45$, $p<0.05$). Post-hoc Tukey HSD comparisons show that transfers like 1 HP to 2 HP (F1 = 0.997) significantly outperform 1 HP to 3 HP (F1 = 0.986), and 2 HP to 1 HP (F1 = 0.988) surpasses 2 HP to 0 HP (F1 = 0.974). For source loads at 0 HP and 3 HP, no significant differences emerge among cross-load targets ($p\approx 0.08$ and $p\approx 0.06$, respectively). An overall ANOVA across all cross-load combinations confirms significant differences ($F\approx 5.67$, $p<0.01$), with the largest disparity between 1 HP to 2 HP (F1 = 0.997) and 0 HP to 3 HP (F1 = 0.948), which is statistically significant ($p<0.01$).
Table 7 also details the model’s transfer learning performance, assessed via k-fold cross-validation. The model achieves perfect same-domain performance (1.0 for both Bearing Gear $\to$ Bearing Gear and Gearbox Gear $\to$ Gearbox Gear), ensuring high reliability within known domains. Its strong cross-domain performance, notably Gearbox Gear $\to$ Bearing Gear (Accuracy 0.972806), supports scalability across related systems.
4.3.4 Feature Importance Analysis
To elucidate the predictive behavior of the logistic regression model, we analyze its learned coefficients, where positive coefficients enhance the likelihood of the positive class (fault conditions), and negative coefficients align with the negative class (healthy conditions). This analysis, illustrated in Figure 6, reveals features’ relative importance and direction at 0HP, 1Hp, 2 HP and 3 HP load. PSD Mid Freq Ch1 notably (for example in 2 Hp load) exhibits the highest positive coefficient ($\approx+1.9$), underscoring its critical role in fault prediction. Other influential positive features include PSD High Freq Ch1 and FFT High Freq Ch2, while features like FFT Mean Ch2, Skew Ch1, and RMS Ch2 display strong negative coefficients, likely reflecting baseline or healthy states.
Figure 6: Logistic regression coefficients showing the relative importance and direction of each feature for 0HP, 1 HP, 2 HP and 3 Hp load.
Here, Ch1 and Ch2 denote Channel 1 and 2, respectively, representing distinct vibration signal acquisition channels from the monitored system. The analysis indicates that Channel 1 features are more predictive of faults, whereas Channel 2 features correlate with regular operation. Frequency-domain features, such as PSD and FFT, dominate across both channels, with time-domain statistical measures (e.g., skewness, RMS) also contributing, albeit with channel-specific variations.
Graph-theoretic properties, including Modularity (Q), Spectral Gap, and Average Path Length, exhibit lower but non-negligible coefficients. While not primary drivers in the linear model, these features capture structural signal characteristics that aid classification.
To further investigate the contribution of features in the SU gearbox and bearing datasets, Figure 7 presents a heatmap of learned coefficients for the gearset/bearing data. This visualization highlights the influence of time- and frequency-domain features across various sensor channels. Warm colors (red) indicate a positive contribution to fault classification, while cool colors (blue) represent negative contributions. In the gearbox dataset, frequency-domain features such as FFT High Freq and FFT Low Freq in Channels 4 and 5 show strong positive influence, underscoring their importance. Time-domain statistical features like Mean, Median, and Kurtosis from the earlier channels also demonstrate notable relevance. Conversely, Channels 6 to 8 exhibit mostly neutral weights, indicating a relatively lower impact. A similar heatmap can be generated for the bearing dataset to analyze the influence of different features on classification decisions
Figure 7: Heatmap of logistic regression coefficients for the SU gearset and bearing dataset, illustrating the contribution of time- and frequency-domain features across sensor channels. Red shades indicate a positive influence on classification, while blue shades denote negative contributions.
This analysis highlights the interpretability of logistic regression, guiding effective feature selection for fault classification. Extending this approach to other datasets can ensure consistency and generalizability across diverse operating conditions.
4.3.5 Sensitivity Analysis
Figure 8: Classification accuracy under increasing levels of feature noise across different operating conditions. The top row shows performance on the CWRU bearing dataset under four machine load settings (0 HP, 1 HP, 2 HP, and 3 HP). The bottom row shows results on the SU Bearing and SU Gearbox datasets. As noise increases, accuracy declines, with higher loads exhibiting greater sensitivity to noise. In contrast, the SU datasets demonstrate stronger robustness across all noise levels.
This study further investigates the robustness of machine learning models to feature noise using three distinct vibration datasets: the CWRU bearing dataset, evaluated under varying machine loads (0 HP to 3 HP), and the SU Bearing and SU Gearbox datasets, assessed under standard operating conditions, as illustrated in Figure 8. In each case, Gaussian noise with increasing standard deviations (0, 0.05, 0.1, 0.2, and 0.5) is added to the input features, and its impact on classification accuracy is evaluated through k-fold cross-validation.
For the CWRU dataset, model performance remains stable under light noise at all load levels, but accuracy steadily declines as noise increases. At 0 HP, the model maintains high accuracy (approximately 0.982) up to a noise level of 0.2, after which it drops moderately to around 0.955 at 0.5 noise. As the load increases to 1 HP and 2 HP, the model becomes more sensitive, with accuracy decreasing to around 0.89 and 0.86, respectively, at the highest noise level. The most significant degradation is observed at 3 HP, where accuracy falls from approximately 0.982 to below 0.88 under maximum noise, indicating that higher operational loads exacerbate the model’s sensitivity to noisy input features.
In contrast, the SU datasets exhibit strong robustness to noise. The model starts with very high baseline accuracy (approximately 0.992 and 0.994, respectively) and maintains excellent performance up to moderate noise levels (0.2 standard deviation) for both the SU Bearing and SU Gearbox datasets. Even at the highest noise level (0.5), accuracy only dips slightly—to around 0.982 for SU Bearing and 0.979 for SU Gearbox—highlighting their stability and noise resilience. These findings suggest that while noise has a more pronounced effect under varying load conditions in the CWRU dataset, the models trained on the SU datasets remain consistently robust, making them highly reliable for real-world fault diagnosis, even in noisy signals.
4.4 Comparative Analysis
The proposed bearing fault detection model outperforms deep learning approaches (e.g., CNN, LSTM) and graph-based methods (e.g., GCN, GAT) on the CWRU and SU datasets. By integrating feature engineering, signal processing, and graph metrics (e.g., path length $L$, modularity $Q$, spectral gap) with logistic regression, the model achieves superior accuracy, computational efficiency, and interpretability. Performance benchmarks in Table 8 (accuracy) and Table 9 (inference time) demonstrate its clear advantage over more complex alternatives.
Table 8: Comparative Performance of Bearing Fault Detection Models on CWRU and PU Datasets
Method CWRU Accuracy (%) SU Accuracy (%)
MBH-LPQ-VGGish (Chennana et al., 2025) 98.95
VGG-16 (Chennana et al., 2025) 96.84
YamNet (Chennana et al., 2025) 71.52
VGGish (Chennana et al., 2025) 98.42
MBH-LPQ (Chennana et al., 2025) 97.89
DenseNet-12 (fine-tuned) (Niyongabo et al., 2022) 98.57
Stacking Classifier (Makrouf et al., 2023) 97.20
TDANET (Li et al., 2024) 97.69
WDCNN (Gao et al., 2024) 84.57
MFFCNN (Gao et al., 2024) 86.47
GAE (Singh, 2025) 97.6
Proposed Model 99.89 99.99
SCNN 1D (Xu et al., 2025) 99.60
SCNN 2D (Xu et al., 2025) 99.67
SCNN 3D (Xu et al., 2025) 99.93
Inception-LSTM (Wei et al., 2024) 99.50
Sparse Transformer (Zhou et al., 2024) 99.98
Table 9: Inference Time Comparison on the CWRU and SU Dataset
Model CWRU Inference Time (s) SU Inference Time (s)
CBAM-MFFCNN (Gao et al., 2024) 0.1179
MBH-LPQ (Chennana et al., 2025) 0.224
VGGish (Chennana et al., 2025) 1.769
VGGish+MBH-LPQ (Chennana et al., 2025) 1.998
OCSSA-VMD-CNN-BiLSTM (Chang and Bao, 2024) 8.470
PSO-tuned XGBoost (Lee and Maceren, 2024) 127.212
DTL-Res2Net-CBAM (Wang and Zhang, 2024) 372.000
Proposed Model 0.000252 0.0001639
SCNN (Xu et al., 2025) 0.002
Shuffle Net (Xu et al., 2025) 0.060
Mobile Net (Xu et al., 2025) 0.080
DL-CNN (Shao et al., 2018) 536
The model transforms vibrational data into a $(20ch+3)$-dimensional feature vector, where $C$ denotes the number of channels. For each channel, it extracts 8 time-domain features (e.g., mean, RMS, skewness), 12 frequency-domain features (e.g., dominant FFT frequency, PSD peak), and 3 graph-based metrics, resulting in 43 features when $ch=2$ (as in the CWRU and GearBox datasets). Features such as FFT Dom Freq Ch1 capture fault harmonics (e.g., 50–100 Hz on CWRU, 100–150 Hz on SU), while graph metrics like $Q$ quantify structural clustering. When trained and tested on the same dataset without cross-dataset learning, logistic regression achieves 99.89% accuracy on CWRU and 100.00% on SU (Table 8), outperforming more complex deep learning models and graph convolutional networks, which often suffer from heavy preprocessing requirements and limited interpretability.
On CWRU, the model’s 99.89% accuracy is compatible with MBH-LPQ-VGGish (98.95%) (Chennana et al., 2025), out performs DenseNet-12 (98.57%) (Niyongabo et al., 2022), and WDCNN (84.57%) (Gao et al., 2024). On SU, it achieves 100.00%, surpassing SCNN 3D (99.93%) (Xu et al., 2025) and Sparse Transformer (99.98%) (Zhou et al., 2024).
The model’s inference time is 0.000252 s on CWRU and 0.0001639 s on SU (Table 9), significantly faster than CBAM-MFFCNN (0.1179 s) (Gao et al., 2024), VGGish (1.769 s) (Chennana et al., 2025), MBH-LPQ-VGGish (1.998 s)  (Chennana et al., 2025), SCNN (0.002 s) (Xu et al., 2025), and DL-CNN (536 s) (Shao et al., 2018). Precomputed features reduce complexity, enabling real-time fault detection, unlike resource-intensive deep learning and GCNs.
SHAP analysis enhances the model’s interpretability. For eaxmple, the feature FFT Dom Freq Ch1 (SHAP: -15 to +15 on SU gearbox dataset) links to fault frequencies, and $Q$ (SHAP: 2–4) captures clustering. Similar trends appear on CWRU (SHAP: -20 to +20). Unlike deep learning and GCNs, which rely on vague post-hoc methods like Grad-CAM, the model’s features directly connect to physical phenomena, simplifying diagnosis.
The model performs well with limited data, achieving 99.89% on CWRU , 99.99% on SU with small subgraphs ($|\mathcal{G}_{i}|\leq 200$). Engineered features ensure strong generalization, unlike LSTMs (F1: 0.90 on small data) or GCNs, which falter without large graph samples.
In transfer learning experiments, the direction of training plays a critical role in determining generalization performance. On the CWRU dataset, training the model on 1 HP load data and testing across other load conditions (0, 2, and 3 HP) yields a higher average accuracy (99.00%) than training on 3 HP data (98.63%). This indicates that features learned under moderate loading conditions (1 HP) are more transferable, likely because they encapsulate characteristics of both low and high load dynamics. A similar trend is observed on the SU dataset, where training on Gearbox Gear data results in better transfer accuracy (97.3%) when tested on Bearing Gear data, compared to the reverse direction (94.9%). Notably, existing methods do not perform true transfer learning; instead, they train and test on mixed data from each vibrational dataset. This practice can lead to inflated performance metrics and limits the evaluation of a model’s ability to generalize across different domains or operating conditions (Zhao et al., 2019).
4.5 Discussion
The results from both the CWRU and SU datasets demonstrate the effectiveness and adaptability of the proposed graph-based fault detection approach across a wide range of operational scenarios and fault conditions. One of the most significant observations is the model’s strong performance across varying machine loads, particularly in the CWRU dataset. While accuracy and F1-scores were consistently high across all load conditions, classification performance slightly dipped under low-load scenarios (e.g., 0 HP), where vibration signals are less distinguishable due to weaker fault-induced signatures. In contrast, higher loads (2 HP and 3 HP) produced more pronounced fault features, contributing to improved classification metrics and better graph separability.
In the SU datasets, where operational settings were more controlled and balanced across fault types, the model demonstrated nearly perfect classification across all fault classes. These results highlight the robustness of the proposed method in identifying subtle vibration pattern changes, even across datasets with differing domain characteristics.
Additionally, the transfer learning experiments revealed that while same-load source–target configurations led to near-perfect classification, cross-load performance remained strong but showed slight drops depending on the disparity in load conditions. Statistical analysis validated that these variations were significant, reinforcing the need for adaptive models or fine-tuning in real-world, variable-load deployments.
Another key finding relates to graph construction parameters. Adjusting $k_{max}$ and subgraph sizes significantly influenced model efficiency and structural clarity. Small-world characteristics and higher modularity scores were more pronounced in mid-sized subgraphs, balancing detection performance with computational efficiency. Feature noise analysis further confirmed the model’s resilience, especially in the SU datasets, where accuracy remained stable even under high noise levels.
5 Conclusion
This paper presents a novel, entropy-guided, graph-based framework for multiclass fault detection in rotating machinery, tested across two benchmark datasets (CWRU and SU). By optimizing signal segmentation, extracting time–frequency features, and encoding them into graph structures, the proposed method enables interpretable and high-precision fault classification. Graph-theoretic metrics such as average path length, modularity, and spectral gap provide insights into structural shifts caused by different fault types, enhancing model discriminability and robustness.
Extensive experiments demonstrate the model’s adaptability across diverse operating conditions, load levels, and fault categories. It exhibits near-perfect classification accuracy on both bearing and gearbox datasets from SU and delivers consistently strong performance under varying loads in the CWRU dataset. Furthermore, transfer learning experiments confirm the model’s ability to generalize across different operational settings, with only slight performance drops in mismatched load scenarios.
The approach also proves resilient to feature noise, particularly in SU datasets, highlighting its potential for real-world deployments. Future work may extend this framework to more complex, multivariate signal scenarios and explore the integration of domain adaptation techniques to further improve cross-condition generalization.
Statements and Declarations
Competing Interests
The authors declare that there are no competing interests associated with this research work.
Funding
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Informed Consent
Informed consent was obtained from all individual participants included in the study.
Data Availability
The datasets generated and/or analyzed during the current study are available in upon reasonable request from the corresponding author.
Declaration of AI Assistance
The authors acknowledge the use of AI-assisted tools, including ProWritingAid, during the preparation of this article. These tools were employed solely for improving grammar, sentence structure, and language refinement. The research, analysis, and conclusions drawn in this article are entirely the authors’ original work, and no AI tool was used to generate intellectual content or conduct the underlying research.